# Spatial Transformers for Radio Map Estimation

### v1

@@ IV-A Feature Design
Maxwell’s equations dictate that RME exhibits certain invariance properties, such as translation and rotation invariance. Note that if the estimation of a map as a whole is considered rather than the estimation of a map at a single location ${\color[rgb]{0,0,0}\bm{r}}$, these invariances become equivariances. If the invariances of an problem are not imposed by the estimator architecture, they must be learned from data, which increases drastically the amount of training data required to attain a target performance. This motivates a feature design that enforces these invariances.
@@ IV-B Dataset Preparation
The data to train and test an RME estimator typically consists of a collection of measurements collected in multiple geographical areas along with their locations. Using this data, one needs to generate sets of training and testing examples. Since the proposed estimator is gridless, the procedure differs from the one used in existing DNN estimators.
In particular, one can proceed as follows to generate the ${\color[rgb]{0,0,0}t}$-th example. First, select one of the geographical areas. Among the measurements in this area, choose one as the target and ${{\color[rgb]{0,0,0}N}}[{\color[rgb]{0,0,0}t}]$ of them as the input. As discussed later, the value of ${{\color[rgb]{0,0,0}N}}[{\color[rgb]{0,0,0}t}]$ must be selected depending on the training or testing goals. With the ${{\color[rgb]{0,0,0}N}}[{\color[rgb]{0,0,0}t}]$ measurements, their locations, and the location of the target measurement, construct the feature matrix ${\color[rgb]{0,0,0}\bm{X}}[{\color[rgb]{0,0,0}t}]$ as indicated in Sec. IV-A. By denoting the target measurement as ${\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}[{\color[rgb]{0,0,0}t}]$, the dataset can be expressed as $\left\{({\color[rgb]{0,0,0}\bm{X}}[{\color[rgb]{0,0,0}t}],{\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}[{\color[rgb]{0,0,0}t}])\right\}_{{\color[rgb]{0,0,0}t}=1}^{{\color[rgb]{0,0,0}T}}$.
@@ IV-C Architecture
The proposed estimator adopts a transformer architecture and it comprises the blocks in the shaded area of Fig. 2. The rest of the blocks are used for active sensing and are described later. The feature vectors ${\color[rgb]{0,0,0}\bm{x}}_{{{\color[rgb]{0,0,0}n}}}$ are first passed separately through a linear layer $L$ that increases their dimension to ${\color[rgb]{0,0,0}D}$. Then, a composition of attention blocks is applied. The input and output of every block are vectors of dimension ${\color[rgb]{0,0,0}D}$. Finally, the output vectors of the last attention block are passed through a linear layer that reduces their dimension to 1. The returned ${{\color[rgb]{0,0,0}N}}$ scalars can be collected in the vector ${\color[rgb]{0,0,0}F}({\color[rgb]{0,0,0}\bm{X}})\triangleq[{\color[rgb]{0,0,0}F}_{1}({\color[rgb]{0,0,0}\bm{X}}),\ldots,{\color[rgb]{0,0,0}F}_{{\color[rgb]{0,0,0}N}}({\color[rgb]{0,0,0}\bm{X}})]^{\top}$. To obtain an estimate ${{\color[rgb]{0,0,0}\hat{{\color[rgb]{0,0,0}\gamma}}}}({\color[rgb]{0,0,0}\bm{r}})$, these ${{\color[rgb]{0,0,0}N}}$ scalars can be reduced into a single one, e.g. by averaging. One can then train by minimizing the mean square error:
@@
The limitation of this approach is that examples for all the necessary values of ${{\color[rgb]{0,0,0}N}}[{\color[rgb]{0,0,0}t}]$ need to be included in the dataset. This may result in an unnecessarily very large dataset. To alleviate this issue, causal self-attention is commonly used in the context of transformers. As discussed later, it is not fully suitable for RME, but the training time reduction may pay off. A causal self-attention head ${\color[rgb]{0,0,0}H}_{\text{c}}$ is similar to the self-attention head introduced in Sec. III-A, but the ${\color[rgb]{0,0,0}n}$-th output vector is only allowed to depend on the input vectors $\{{\color[rgb]{0,0,0}\bm{x}}_{{{\color[rgb]{0,0,0}n}}^{\prime}}\}_{{{\color[rgb]{0,0,0}n}}^{\prime}=1}^{{\color[rgb]{0,0,0}n}}$. In other words, the ${\color[rgb]{0,0,0}n}$-th column of the matrix ${\color[rgb]{0,0,0}H}_{\text{c}}({\color[rgb]{0,0,0}\bm{X}})$ is ${\color[rgb]{0,0,0}H}({\color[rgb]{0,0,0}\bm{X}}_{{\color[rgb]{0,0,0}n}},{\color[rgb]{0,0,0}\bm{x}}_{{\color[rgb]{0,0,0}n}})$, where ${\color[rgb]{0,0,0}\bm{X}}_{{\color[rgb]{0,0,0}n}}\triangleq[{\color[rgb]{0,0,0}\bm{x}}_{1},\ldots,{\color[rgb]{0,0,0}\bm{x}}_{{\color[rgb]{0,0,0}n}}]$. An attention block that uses a causal self-attention head will be denoted as ${\color[rgb]{0,0,0}B_{\text{c}}}$.
@@
where now ${{\color[rgb]{0,0,0}N}}[{\color[rgb]{0,0,0}t}]={{\color[rgb]{0,0,0}N}}$ for all ${\color[rgb]{0,0,0}t}$. The main limitation of this approach is the loss of the invariance to the permutation of the measurements. This is not a problem in NLP since the tokens in a word are ordered, but it is a problem in RME since the measurements are not ordered. However, this is the price to be paid for the reduction in training time.
@@ V Attention-based Active Sensing
To this end, process $\{{\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}}\}_{{{\color[rgb]{0,0,0}n}}={{\color[rgb]{0,0,0}N}}+1}^{{{\color[rgb]{0,0,0}N}}+{\color[rgb]{0,0,0}M}}$ like the measurement locations, with the same translation and rotation as used to obtain ${\color[rgb]{0,0,0}\bm{X}}$. This yields the matrix ${\color[rgb]{0,0,0}{\bar{\bm{X}}}}$, which has one row less than ${\color[rgb]{0,0,0}\bm{X}}$ since it does not contain the measurement values. Decomposing ${\color[rgb]{0,0,0}F}$ into an encoder and a decoder part ${\color[rgb]{0,0,0}F}^{\text{D}}$ and ${\color[rgb]{0,0,0}F}^{\text{E}}$, the architecture of STORM is extended to obtain:
@@
To train STORM to predict the quality of each candidate estimate without using the measurements at the candidate locations, the idea here is to construct a combined estimate as a convex combination of the candidate estimates. The weights in this convex combination are the entries of ${\color[rgb]{0,0,0}C}({\color[rgb]{0,0,0}\bm{X}},{\color[rgb]{0,0,0}{\bar{\bm{X}}}})$. The loss becomes:
@@ VI Experiments with Synthetic and Real Data
This section evaluates the performance of STORM on three datasets. When it comes to map estimation error, STORM is compared with three non-DNN and four DNN estimators. The non-DNN estimators include K-nearest neighbors (K-NN), Kriging [7, 8], and kernel ridge regression (KRR); see [6] and references therein. These estimators are trained as described in [22]. The compared DNN estimators include  [DNN 1] the completion autoencoder in [15],  [DNN 2] the U-Net from [14],  [DNN 3] the U-net from [13],  and [DNN 4] the autoencoder in [8]. Unless stated otherwise, STORM uses multi-head attention with 2 heads and embedding dimension ${\color[rgb]{0,0,0}D}=48$, which results in around 100 k parameters.
Each of the considered datasets consists of several measurement sets (MSs). Each one is a set of measurements collected in an ${\color[rgb]{0,0,0}\underline{L}}_{\text{x}}\times{\color[rgb]{0,0,0}\underline{L}}_{\text{y}}$ rectangular environment. Each MS corresponds to a different function ${{\color[rgb]{0,0,0}\gamma}}$. These MSs are split into training and testing MSs. The training measurements are obtained by selecting the measurements inside randomly drawn ${\color[rgb]{0,0,0}L}\times{\color[rgb]{0,0,0}L}$ square patches inside a randomly selected MS. Likewise for the testing measurements. To favor the competing grid-aware DNN estimators, the aforementioned square patches are aligned with the grid in which the measurements are collected. Worse performance of these estimators is expected otherwise. The grid spacing is denoted as ${\color[rgb]{0,0,0}\Delta}$.
@@
The first dataset in this paper is generated by Remcom’s Wireless InSite ray-tracing software using a 3D model of an area in downtown Rosslyn, Virginia, with ${\color[rgb]{0,0,0}\underline{L}}_{\text{x}}\approx{\color[rgb]{0,0,0}\underline{L}}_{\text{y}}\approx 700$ m. Each MS corresponds to a different transmitter location. Fig. 3 shows the RMSE vs. ${{\color[rgb]{0,0,0}{{\color[rgb]{0,0,0}N}}}}$.
@@
Figure 5: RMSE for 4G gradiant data vs. ${{\color[rgb]{0,0,0}{{\color[rgb]{0,0,0}N}}}}$ when ${\color[rgb]{0,0,0}L}=64$ m, ${\color[rgb]{0,0,0}\Delta}=4$ m, and the estimators are trained with ${{\color[rgb]{0,0,0}{{\color[rgb]{0,0,0}N}}}}\in[20,100]$.
@@
Figure 6: RMSE vs. the number of observations for the active sensing problem with ray-tracing and USRP data. ${\color[rgb]{0,0,0}D}=20$.

### v2

@@ IV-A Feature Design
Maxwell’s equations dictate that RME exhibits certain invariance properties, such as translation and rotation invariance. Note that if the estimation of a map as a whole is considered rather than the estimation of a map at a single location ${\color[rgb]{0,0,0}\bm{r}}$, these invariances become equivariances. If the invariances of a problem are not imposed by the estimator architecture, they must be learned from data, which increases drastically the amount of training data required to attain a target performance. This motivates a feature design that enforces these invariances.
@@
The exp yields positive weights if ${\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}_{{{\color[rgb]{0,0,0}n}}}$ is in dB units.
@@ IV-B Dataset Preparation
The data to train and test an RME estimator typically consists of a collection of measurement sets (MSs). Each MS contains geolocated measurements corresponding to a different true ${\color[rgb]{0,0,0}\gamma}$. For example, each MS can be collected in a different geographical area or for different transmitter locations. MSs are then used to generate sets of training and testing examples. Since the proposed estimator is gridless, the procedure differs from the one used in existing DNN estimators.
In particular, one can proceed as follows to generate the ${\color[rgb]{0,0,0}t}$-th example. First, randomly select one of the MSs. Among the measurements in MS, choose one as the target and ${{\color[rgb]{0,0,0}N}}[{\color[rgb]{0,0,0}t}]$ of them as the input. As discussed later, the value of ${{\color[rgb]{0,0,0}N}}[{\color[rgb]{0,0,0}t}]$ must be selected depending on the training or testing goals. With the ${{\color[rgb]{0,0,0}N}}[{\color[rgb]{0,0,0}t}]$ measurements, their locations, and the location of the target measurement, construct the feature matrix ${\color[rgb]{0,0,0}\bm{X}}[{\color[rgb]{0,0,0}t}]$ as indicated in Sec. IV-A. By denoting the target measurement as ${\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}[{\color[rgb]{0,0,0}t}]$, the dataset can be expressed as $\left\{({\color[rgb]{0,0,0}\bm{X}}[{\color[rgb]{0,0,0}t}],{\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}[{\color[rgb]{0,0,0}t}])\right\}_{{\color[rgb]{0,0,0}t}=1}^{{\color[rgb]{0,0,0}T}}$.
@@ IV-C Architecture
The proposed estimator adopts a transformer architecture and comprises the blocks in the shaded area of Fig. 2. The rest of the blocks are used for active sensing and are described later. The feature vectors ${\color[rgb]{0,0,0}\bm{x}}_{{{\color[rgb]{0,0,0}n}}}$ are first passed separately through a linear layer $L$ that increases their dimension to ${\color[rgb]{0,0,0}D}$. Then, a composition of attention blocks is applied. The input and output of every block are vectors of dimension ${\color[rgb]{0,0,0}D}$. Finally, the output vectors of the last attention block are passed through a linear layer that reduces their dimension to 1. The returned ${{\color[rgb]{0,0,0}N}}$ scalars can be collected in the vector ${\color[rgb]{0,0,0}F}({\color[rgb]{0,0,0}\bm{X}})\triangleq[{\color[rgb]{0,0,0}F}_{1}({\color[rgb]{0,0,0}\bm{X}}),\ldots,{\color[rgb]{0,0,0}F}_{{\color[rgb]{0,0,0}N}}({\color[rgb]{0,0,0}\bm{X}})]^{\top}$. To obtain an estimate ${{\color[rgb]{0,0,0}\hat{{\color[rgb]{0,0,0}\gamma}}}}({\color[rgb]{0,0,0}\bm{r}})$, these ${{\color[rgb]{0,0,0}N}}$ scalars could be reduced into a single one, e.g. by averaging, and train by minimizing the mean square error:
@@
The limitation of this approach is that examples for all the necessary values of ${{\color[rgb]{0,0,0}N}}[{\color[rgb]{0,0,0}t}]$ need to be included in the dataset. This may result in an unnecessarily large dataset. To alleviate this issue, causal self-attention is commonly used in the context of transformers. As discussed later, it is not fully suitable for RME, but the training time reduction may pay off. A causal self-attention head ${\color[rgb]{0,0,0}H}_{\text{c}}$ is similar to the self-attention head introduced in Sec. III-A, but the ${\color[rgb]{0,0,0}n}$-th output vector is only allowed to depend on the input vectors $\{{\color[rgb]{0,0,0}\bm{x}}_{{{\color[rgb]{0,0,0}n}}^{\prime}}\}_{{{\color[rgb]{0,0,0}n}}^{\prime}=1}^{{\color[rgb]{0,0,0}n}}$. In other words, the ${\color[rgb]{0,0,0}n}$-th column of the matrix ${\color[rgb]{0,0,0}H}_{\text{c}}({\color[rgb]{0,0,0}\bm{X}})$ is ${\color[rgb]{0,0,0}H}({\color[rgb]{0,0,0}\bm{X}}_{{\color[rgb]{0,0,0}n}},{\color[rgb]{0,0,0}\bm{x}}_{{\color[rgb]{0,0,0}n}})$, where ${\color[rgb]{0,0,0}\bm{X}}_{{\color[rgb]{0,0,0}n}}\triangleq[{\color[rgb]{0,0,0}\bm{x}}_{1},\ldots,{\color[rgb]{0,0,0}\bm{x}}_{{\color[rgb]{0,0,0}n}}]$. An attention block that uses a causal self-attention head will be denoted as ${\color[rgb]{0,0,0}B_{\text{c}}}$.
@@
where now ${{\color[rgb]{0,0,0}N}}[{\color[rgb]{0,0,0}t}]={{\color[rgb]{0,0,0}N}}$ for all ${\color[rgb]{0,0,0}t}$. The main limitation of this approach is the loss of the invariance to the permutation of the measurements. This is not a problem in NLP since the tokens in a word are ordered, but it is a problem in RME since the measurements are not ordered. However, this is the price to be paid for the reduction in training complexity.
@@ V Attention-based Active Sensing
To this end, process $\{{\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}}\}_{{{\color[rgb]{0,0,0}n}}={{\color[rgb]{0,0,0}N}}+1}^{{{\color[rgb]{0,0,0}N}}+{\color[rgb]{0,0,0}M}}$ identically to the measurement locations, with the same translation and rotation as used to obtain ${\color[rgb]{0,0,0}\bm{X}}$. This yields the matrix ${\color[rgb]{0,0,0}{\bar{\bm{X}}}}$, which has one row less than ${\color[rgb]{0,0,0}\bm{X}}$ since it does not contain the measurement values. Decomposing ${\color[rgb]{0,0,0}F}$ into an encoder part ${\color[rgb]{0,0,0}F}^{\text{E}}$ and a decoder part ${\color[rgb]{0,0,0}F}^{\text{D}}$, the architecture of STORM is extended to obtain:
@@
To train STORM to predict the quality of each candidate estimate without using the measurements at the candidate locations, the idea here is to construct a combined estimate as a convex combination of the candidate estimates. The weights in this convex combination are the entries of ${\color[rgb]{0,0,0}C}({\color[rgb]{0,0,0}\bm{X}},{\color[rgb]{0,0,0}{\bar{\bm{X}}}})$, which do not depend on those measurements. The loss becomes:
@@ VI Experiments with Synthetic and Real Data
This section evaluates the performance of STORM on three datasets. When it comes to map estimation error, STORM is compared with three non-DNN and four DNN estimators. The non-DNN estimators include K-nearest neighbors (KNN), Kriging [7, 8], and kernel ridge regression (KRR); see [6] and references therein. These estimators are trained as described in [22]. The compared DNN estimators include  [DNN 1] the completion autoencoder in [15],  [DNN 2] the U-Net from [14],  [DNN 3] the U-net from [13],  and [DNN 4] the autoencoder in [8]. Unless stated otherwise, STORM uses multi-head attention with 2 heads and embedding dimension ${\color[rgb]{0,0,0}D}=48$, which results in around 100 k parameters.
Each of the considered datasets comprises multiple MSs collected in an ${\color[rgb]{0,0,0}\underline{L}}_{\text{x}}\times{\color[rgb]{0,0,0}\underline{L}}_{\text{y}}$ rectangular environment. These MSs are split into training and testing MSs. To obtain each training example, one can collect the measurements inside an ${\color[rgb]{0,0,0}L}\times{\color[rgb]{0,0,0}L}$ square patch drawn uniformly at random and included in a training MS selected uniformly at random. Likewise for the testing examples. To favor the competing (grid-aware) DNN estimators, the aforementioned square patches are aligned with the grid in which the measurements are collected. Worse performance of these estimators is expected otherwise. The grid spacing is denoted as ${\color[rgb]{0,0,0}\Delta}$.
@@
The first dataset in this paper is generated by Remcom’s Wireless InSite ray-tracing software using a 3D model of an area in downtown Rosslyn, Virginia, with ${\color[rgb]{0,0,0}\underline{L}}_{\text{x}}\approx{\color[rgb]{0,0,0}\underline{L}}_{\text{y}}\approx 700$ m. Each MS corresponds to a different transmitter location; see [15] for details. Fig. 3 shows the RMSE vs. ${{\color[rgb]{0,0,0}{{\color[rgb]{0,0,0}N}}}}$.
@@
Figure 5: RMSE for 4G data vs. ${{\color[rgb]{0,0,0}{{\color[rgb]{0,0,0}N}}}}$ when ${\color[rgb]{0,0,0}L}=64$ m, ${\color[rgb]{0,0,0}\Delta}=4$ m, and the estimators are trained with ${{\color[rgb]{0,0,0}{{\color[rgb]{0,0,0}N}}}}\in[20,100]$.
@@
Figure 6: RMSE vs. ${{\color[rgb]{0,0,0}{{\color[rgb]{0,0,0}N}}}}$ for the active sensing problem with ray-tracing and USRP data. ${\color[rgb]{0,0,0}D}=20$.